EXERCISE 6.4. In the proof of the Schröder-Bernstein Theorem, define a function 19-(x) if TEX GC) = f(x) if TEX (5(x) if X.
Answer #1

As per the notations in the proof of the Schroder-Bernstein Theorem, Xe,Xo and Xi are the sets of all those elements of X which have an even or odd or infinite number of ancestors, respectively. Similarly, for Y.

Now, {Xe,Xo,Xi} and {Ye,Yo,Yi} are partitions of X and Y respectively.

I. The restriction f|​​​​​​Xe : Xe\rightarrow Yo as given x in Xe, f(x) is in Yo and since x has an even number of predecessors g-1(x),f-1(g-1(x)),g-1(f-1(g-1(x)))........f-1(g-1(....g-1(x)..)), hence f(x) clearly has an odd number of predecessors (namely x = f-1(f(x)) and all the above in the previous list).

Now, since f is injective, hence the above restriction is injective.

Further , given y in Yo, then y has an odd number of predecessors and hence, y has at least one predecessor and hence the immediate predecessor f-1(y). Obviously f-1(y) has an even number of predecessors and hence f-1(y) is in Xe. Further, f(f-1(y)) = y. Thus, f|Xe is surjective. Thus, f|Xe : Xe\rightarrow Yo is a bijection.

II. Consider g-1 : g(Y) \rightarrow Y. Since, x is in Xi U Xo implies that x has at least one predecessor and hence, an immediate predecessor, then g-1(x) exists as an element in Y. In other words, x = g(g-1(x)) is in g(Y). Thus, Xi U Xo is a subset of g(Y).

Consider the restriction g-1|Xo : Xo\rightarrow Y. Since each x in Xo has an odd number of predecessors, hence g-1(x) has an even number of predecessors(namely all the predecessors of x except g-1(x) itself). Thus, g-1|Xo : Xo\rightarrow Ye. Since g is injective, hence g-1 is injective. Thus, the restriction g-1|Xo is injective. Now, let y be in Ye. Then, g(y) has an odd number of predecessors(namely y = g-1(g(y)) and all the evenly many predecessors of y). Thus g(y) is in Xo and further, g-1(g(y)) = y. Thus, g-1|Xo is surjective. Hence, g-1|Xo : Xo\rightarrow Ye is bijective.

Again, consider the restriction g-1|Xi : Xi\rightarrow Y. If x is in Xi, then x has infinitely many predecessors g-1(x),f-1(g-1(x), ...... Thus, g-1(x) has infinitely many predecessors(all predecessors of x except g-1(x) itself). Thus, g-1|Xi : Xi\rightarrow Yi. Again, since g is injective, so is g-1|Xi. Finally, given y in Yi, since y has infinitely many predecessors, hence g(y) has infinitely many predecessors (namely, y = g-1(g(y)) and all the predecessors of y). Thus, g​​​​​​(y) is in Xi and g-1(g(y)) = y. Thus, g-1|Xi is surjective. Hence, g-1|Xi : Xi\rightarrow Yi is bijective.

Now, stitching all the bijective maps f|Xe : Xe \rightarrow Yo, g-1|Xo : Xo\rightarrow Ye and g-1|Xi : Xi\rightarrow Yi together and using the fact that {Xe,Xi,Xo} and {Ye,Yi,Yo} are partitions of X and Y respectively, we observe that the map

G : X \rightarrow Y given by,

G(x) = { g-1(x) if x is in Xi

f(x) if x is in Xe

g-1(x) if x is in Xo

is a bijection.

Here is a picture representing this proof:

*Note that f maps Xe onto Yo, g-1 maps Xo onto Ye and g-1 maps Xi onto Yi.

a dud be ?, 72.07. it of 66 Xi x ex Sway hogy